Solve the following equations.   x/3 + 1 = 7/15 
Solve the following equations.   x/3 + 1 = 7/15 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x/3+1=7/15
x/3=7/15–1
x/3 = (7 -15)/15
x/3 = -8/15
x = -8/15 × 3
x=-8/5

Thanks